# Tunability of the magnetic properties in Ni intercalated transition metal dichalcogenide NbSe2

### v1

@@ II.4 Nearest-neighbor spin exchange coupling
Here, represents the total spin vector of the Ni atoms. For NiNb4Se8, Ni exists in an electronic configuration close to the Ni3+(3d7) state. This means that the triply degenerate t2g states are fully occupied, while the doubly degenerate eg states accommodate a single electron. Consequently, Ni adopts a low-spin state with a total spin of =. The magnetic moments extracted from the ground state phase AFM2 range from 0.436 at U=0 to 0.975 at U=3 eV, in agreement with an = electronic state.
@@
While is the nonmagnetic total energy of the system. The same equations are valid without and with SOC. To determine the nearest-neighbor exchange parameter from DFT calculations, we employ the Heisenberg Hamiltonian and establish a relationship between the total energies of just three distinct magnetic configurations: From the previous set of equations, we will consider the first three equations (2-4) to derive J∥ and J⟂. These equations will be used in the next Section to derive the magnetic exchanges J∥ and J⟂ in the non-relativistic and relativistic cases.
@@ III Electronic and magnetic properties with SOC: role of the MCA
In this section, we report the study of the magnetic phases as a function of the electronic correlations for NiNb4Se8 in the relativistic case. With respect to the previous section, we include the investigation of the magnetocrystalline anisotropy and the study of the non-collinear 120∘ phase. Indeed, given the triangular lattice in the plane, if the system is antiferromagnetic in the plane, the magnetic system could become frustrated and the magnetism will become noncollinearDey and Botana (2020).
@@ III.2 Magnetic ground state with SOC: NN magnetic couplings
Using the same strategy used in the previous section, we have calculated the energy differences of the four collinear magnetic phases in the presence of SOC. The results are reported in Fig. 7. Comparing these results with the ones without SOC, we obtain a reduction of UC to a value around 1.1 eV. For the rest, the calculation results reveal that SOC induces minor changes in the ground state energies of the four magnetic systems, though the overall trend remains largely unchanged. FM and AFM2 remain the most stable configurations. As in the case without SOC, the energy difference between the AFM1 and NM states is nearly identical at U=0, likely due to the small magnetic moment (0.27 with SOC). This will lead to a small compared to the results for other values of U.
Employing self-consistent energy differences, we calculate the magnetic exchanges and as defined in the previous sections for both relativistic and non-relativistic calculations. The structure used in the calculation was obtained from optimization under the AFM2 magnetic configuration. We compare the value of exchange constants and as a function of U, with and without SOC. Our exchange energy calculations provide a comprehensive explanation for the magnetic ground state energetics. Disregarding the small variation in U and the effects of SOC, the absolute value of is one order of magnitude greater than that of , demonstrating that the out-of-plane exchange energy is nearly significantly higher than the in-plane exchange energy. Therefore, we have a strongly anisotropic magnetism. From the perspective of SOC, it has a slight effect on the in-plane exchange energy , making the ferromagnetic state more favored as the ground state. The significantly negative out-of-plane exchange interactions not only strongly favor FM and AFM2 configurations with identical out-of-plane spin orientations as the dominant ground states, but also reduce the structural stability of the out-of-plane antiparallel spin magnetic configuration. Leading to an increase in the ground state energy of both the AM and AFM1 phases and even a tendency towards the nonmagnetic state. Furthermore, the in-plane exchange interactions are negligibly small, resulting in minimal energy differences between FM and AFM2 configurations with opposite in-plane spin alignments. The influence of U is minimal on , but it significantly affects , even forcing it to change sign, which further alters the magnetic ground state. For low values of U, the positive suggests that the AFM2 phase should be the ground state. However, as U increases, decreases notably, shifting the ground state from AFM2 to FM. Therefore, as previously mentioned, it is challenging to determine whether the ground state is FM or AFM2 based on the existing calculations, as U values between 1 and 2 eV are reasonable for NiNb4Se8.
@@ III.3 Fermiology of the AFM2 and FM phase
We plot the relativistic band structure focusing on the most relevant magnetic phases, which are the FM phase and the AFM2 phase. Their band structures for U=2 eV are plotted in Fig. 8(a) and 8(b), respectively. As in the case without SOC, significant metallicity is found in both cases. In both cases, there are electron bands at the point close to the Fermi level, which is the opposite of the pristine case. The AFM2 phase hosts the Kramers degeneracy.
@@
We plot the total density of states (TDOS) for the FM and AFM2 phases in Fig. 9(a) and 9(b), respectively, since these phases are the ground states in different regions of the phase diagram. In both cases, there is a significant DOS at the Fermi level, which produces the screening of the Coulomb repulsion, lowering the electronic correlations. We compare these DOS results with the pristine NbSe2, in which there is a VHS at -0.2 eV with respect to the Fermi level, particularly due to saddle points in the electronic band structure near the M point in the Brillouin zoneCalandra et al. (2009). The intercalated transition metal d-orbitals bands are narrower than the Nb bands and closer to the Fermi level. Once the Nb and Ni bands hybridize, the Nb bands are pushed away from the Fermi level. The intercalation of the Ni atoms shifts the VHS at -0.4 eV from the Fermi level, as we can see in both Figs. 9(a) and 9(b). As a result, this shift of the VHS reduces the TDOS at the Fermi level and can suppress the electronic instabilities, such as the CDW and superconductivityYan et al. (2019). It was shown that there is a strong interplay between a VHS and a CDW state in the isostructural and isoelectronic 2H-TaSe2Luckin et al. (2024), therefore, we can expect that Ni intercalation would affect the CDW. Indeed, after intercalation, the change of the Fermi surface stabilizes the lattice. On the other hand, Ni electrons are involved in the Nb-Se covalent bonds within each monolayer, which weakens the Nb-Nb interactions and prevents the clustering effect of Nb atoms responsible for the CDW distortion. Thus, because of these two factors, the CDW phase at low temperature is expected to be suppressed in the intercalated NbSe2.Yin et al. (2023).
III.4 Non collinear phases
@@
The magnetic coupling is antiferromagnetic within the plane, therefore, our system is an effective 2D triangular lattice with antiferromagnetic first-neighbor interaction for low U and ferromagnetic first-neighbor interaction for high U. The collinear antiferromagnetic phase investigated in the previous section is the stripe phase. From the literature, we know that the ground state of a 2D triangular system is the stripe when the second neighbor magnetic exchange is antiferromagnetic above a threshold which is around =0.14; while the ground state is non-collinear 120∘ phase when the second neighbor interaction is ferromagnetic or weakly antiferromagneticHu et al. (2015). Between these two phases, the theory predicts the quantum spin liquid phase for insulating systems. Since the system is metallic, it cannot host the quantum spin liquid phase, but the competition between the stripe and the non-collinear phases holds. We constructed a supercell to simulate the non-collinear 120∘ phase, which is reported in Fig. 10. We performed the structural relaxation on the non-collinear phase in the presence of SOC, and we obtained that the non-collinear phase is 15 meV per formula unit higher in energy than the stripe. Therefore, we can assume that the value of is antiferromagnetic and it is the key parameter to stabilize the stripe AFM2 phase. Another factor that assisted the magnetic coupling in stabilizing the stripe phase is the large magnetocrystalline anisotropy. Indeed, the stripe phase can have spins along the easy axis (z-axis), while the non-collinear 120∘ phase has in-plane spins.
@@ IV Conclusions
We performed first-principles calculations for the Ni-intercalated Nb2Se4 with stoichiometry NiNb4Se8. Our results indicate that the magnetic properties of this system are highly tunable, capable of hosting a rich variety of magnetic phases.
NiNb4Se8 is at the border between a ferromagnetic state and a stripe-like antiferromagnetic phase with Kramers degeneracy (referred to as the AFM2 phase). This configuration is characterized by strong out-of-plane ferromagnetic coupling and relatively weak in-plane magnetic interactions. Among the magnetic structures considered, the AFM2 phase emerges as the ground state in the regime of low Coulomb repulsion, consistent with experimental conditions, where screening effects and a deviation from half-filling reduce the effective interaction strength. The system exhibits an easy axis along and a strong in-plane second-nearest-neighbor antiferromagnetic interaction. However, in the case of disordered Ni distribution, the fact that NiNb3S6 has an antiferromagnetic coupling on the triangular lattice in the plane out-of-plane will likely push the experimental samples into a regime of spin-glass. Given the geometrical frustration and proximity to both ferromagnetic and antiferromagnetic regimes, a metamagnetic transition under an applied magnetic field is also plausible. The critical temperature is expected to be below 84 K, which is the Néel temperature of NiNb3S6; a further reduction of the percentage of Ni will push the system into the nonmagnetic phaseWadge et al. (2025).
@@
Ni intercalation significantly alters the electronic structure near the Fermi level and makes the electronic properties more three-dimensional with respect to pristine NbSe2. It shifts the VHS away from the Fermi energy, reducing the TDOS and suppressing electronic instabilities. Nevertheless, the system remains metallic, with a substantial DOS at the Fermi level that provides effective screening of Coulomb interactions. Notable changes are observed in the Fermi surface compared to pristine NbSe2, with multiple sheets appearing in both the FM and AFM2 phases. In the AFM2 and FM phases, there is a switch of the charge carriers at the point since an electron pocket appears at dominated by 3z2-r2 orbital character, with the primary atomic weight localized on the Nb atoms. We look forward to this magnetic system’s experimental realization, which offers a promising platform for exploring tunable magnetism and correlated electronic behavior.

### v2

@@ II.4 Nearest-neighbor spin exchange coupling
Here, represents the total spin vector of the Ni atoms. For Ni0.25NbSe2, Ni exists in an electronic configuration close to the Ni2+(3d8) state. This means that the triply degenerate t2g states are fully occupied, while the doubly degenerate eg states accommodate two electrons. Consequently, Ni adopts a low-spin state with a total spin of =1. The magnetic moments extracted from the ground state phase AFM2 range from 0.436 at U=0 to 0.975 at U=3 eV. Increasing the value of U will give a magnetic moment which is above 1 for the Ni atoms; therefore, the system will adopt a high-spin configuration with the t2g orbitals of the minority channel remaining partially occupied. In the ferromagnetic phase, the Nb atoms exhibit proximitized magnetization, with a magnetic moment of 0.09 at the 2b Wyckoff position and 0.03 at the 6h Wyckoff position for U=2 eV. In all other cases, the magnetization on the Nb atoms is zero due to the compensation arising from the hybridization with the magnetic atoms of opposite spins.
@@
While is the nonmagnetic total energy of the system. The same equations are valid with and without SOC. To determine the nearest-neighbor exchange parameter from DFT calculations, we employ the Heisenberg Hamiltonian and establish a relationship between the total energies of only three distinct magnetic configurations: From the previous set of equations, we will consider the first three equations (2)-(4) to derive J∥ and J⟂. These equations will be used in the next Section to derive the magnetic exchanges J∥ and J⟂ in the non-relativistic and relativistic cases.
II.5 Magnetic coupling at different Ni concentrations
We calculate the magnetic coupling J|| for two other Ni concentrations of NixNbSe2: NiNbSe2 and NiNbSe2, both of which have 2 Ni atoms per unit cell. The lattice constants which were used for Ni0.25NbSe2 were kept the same for different concentrations, so that the Nb-Nb distances are constant for both in-plane and out-of-plane. However, the Ni valence state and the connectivity between the Ni atoms change. In both cases, the system exhibits a ferromagnetic configuration, whereas the configuration with antiferromagnetic coupling breaks time-reversal symmetry, resulting in an altermagnetic phase. We report the band structure of the altermagnetic and ferromagnetic phases of NiNbSe2 in Figs. 8(a) and 8(b), respectively. We report the band structure of the altermagnetic and the ferromagnetic phases of NiNbSe2 in Figs. 9(a) and 9(b), respectively.
Regarding the magnetic ground state, we start from the compound NiNbSe2 and we get the ferromagnetic J||, confirming that J|| is ferromagnetic when the Ni atoms are on top of each other. In the ferromagnetic phase, the system is half metallic and the magnetic moment is 1 . The Ni atoms are in a d9 magnetic configuration, corresponding to an oxidation state of +1. As a consequence, the Nb is in oxidation state +3, making the Nb a 4d2 system similiar to the insulating MoSe2. For this reason, the Nb is depleted from the Fermi level, where only the minority spins of the Ni atoms contribute. Ni would tend to have an oxidation state +2 due to the octahedral crystal field which raises the energy of the 4s level. However, if Ni is +2, Nb would not be able to be +4 as in pristine NbSe2. Therefore, in the stoichiometry NixNbSe2, the Ni atoms have valence +2-, where would tend to increase as the Ni concentration x increases.Boucher et al. (2024) Since for x=1, we have =1, and for x=0 we have =0, and assuming that will depend linearly on the Ni concentration, we can establish that . Therefore, in NixNbSe2 the oxidation state is Ni+2-x, also, the results of the charge at x=1/4 and 1/3 confirms this trend in the first approximation.
We compare the energies of the FM vs AM for the three cases of Ni0.25NbSe2, NiNbSe2 and NiNbSe2. The results are reported in Fig. 10 As previously discussed, the position of Ni atoms in the intercalation layer significantly influences the connectivity. In the case of Ni0.25NbSe2 and NiNbSe2, their identical connectivity leads both systems to favor a FM ground state. In contrast, Ni1/3NbSe2 exhibits a more stable AM configuration, with an energy at least 30 meV/Ni lower than that of the AM states of Ni0.25NbSe2 and NiNbSe2 at U=2 eV. Although not yet experimentally confirmed, our calculations suggest that Ni1/3NbSe2 is a strong candidate material for exhibiting the novel altermagnetic phase.
For the compound NiNbSe2, we have also performed metaGGA calculations to test the stability of the J|| in this class of materials. The employed metaGGA functionals are SCANSun et al. (2015), RSCANBartók and Yates (2019) and R2SCANFurness et al. (2020). We obtain an energy difference EAM-EFM of 32 meV per Ni atom for all the functionals. This result confirms that when the Ni atoms are one on top of each other, the magnetic coupling between them is strongly ferromagnetic. Also in metaGGA, we obtain a half-metallic system for the ferromagnetic phase of NiNbSe2.
@@ III Electronic and magnetic properties with SOC: role of the MCA
In this section, we report the study of the magnetic phases as a function of the electronic correlations for Ni0.25NbSe2 in the relativistic case. Compared to the previous section, we include the investigation of the magnetocrystalline anisotropy and the study of the non-collinear 120∘ phase. Given the triangular lattice in the plane, if the system is antiferromagnetic in the plane, the magnetic system could become frustrated and the magnetism will become noncollinearDey and Botana (2020).
@@ III.2 Magnetic ground state with SOC: NN magnetic couplings
Using the same strategy used in the previous section, we have calculated the energy differences of the four collinear magnetic phases in the presence of SOC. The results are reported in Fig. 11. Comparing these results with the ones without SOC, we obtain a reduction of UC to a value around 1.1 eV. For the rest, the calculation results reveal that SOC induces minor changes in the ground state energies of the four magnetic systems, although the overall trend remains largely unchanged. FM and AFM2 remain the most stable configurations. Similar to the case without SOC, the energy difference between the AFM1 and NM states is nearly identical at U=0, even when SOC is included, likely due to the small magnetic moment (0.27 with SOC). This will lead to a small compared to the results for other values of U.
Employing self-consistent energy differences, we calculate the magnetic exchanges and as defined in the previous sections for both relativistic and non-relativistic calculations. The structure used in the calculation was obtained from optimization under the AFM2 magnetic configuration. We compare the value of exchange constants and as a function of U, with and without SOC. Our exchange energy calculations provide a comprehensive explanation for the magnetic ground state energetics. Disregarding the small variation in U and the effects of SOC, the absolute value of is one order of magnitude greater than that of , demonstrating that the out-of-plane exchange energy is nearly significantly higher than the in-plane exchange energy. Therefore, we have a strongly anisotropic magnetism. From the perspective of SOC, it has a slight effect on the in-plane exchange energy , favoring the ferromagnetic state as the ground state. The significantly negative out-of-plane exchange interactions not only strongly favor FM and AFM2 configurations with identical out-of-plane spin orientations as the dominant ground states, but also reduce the structural stability of the out-of-plane antiparallel spin magnetic configuration. This leads to an increase in the ground state energy of both the AM and AFM1 phases and even a tendency towards the nonmagnetic state. Furthermore, the in-plane exchange interactions are negligibly small, resulting in minimal energy differences between FM and AFM2 configurations with opposite in-plane spin alignments. The influence of U is minimal on , but it significantly affects , even forcing it to change sign, which further alters the magnetic ground state. For low values of U, the positive suggests that the AFM2 phase should be the ground state. However, as U increases, decreases notably, shifting the ground state from AFM2 to FM. Therefore, as previously mentioned, it is challenging to determine whether the ground state is FM or AFM2 based on the existing calculations, as U values between 1 and 2 eV are reasonable for Ni0.25NbSe2.
@@ III.3 Fermiology of the AFM2 and FM phase
We plot the relativistic band structure focusing on the most relevant magnetic phases, which are the FM phase and the AFM2 phase. Their band structures for U=2 eV are plotted in Fig. 12(a) and 12(b), respectively. As in the case without SOC, significant metallicity is found in both cases. In both cases, there are electron bands at the point close to the Fermi level, which is the opposite of the pristine case. The AFM2 phase hosts the Kramers degeneracy.
@@
We plot the total density of states (TDOS) for the FM and AFM2 phases in Fig. 13(a) and 13(b), respectively, since these phases are the ground states in different regions of the phase diagram. In both cases, there is a significant DOS at the Fermi level, which produces the screening of the Coulomb repulsion, lowering the electronic correlations. We compare these DOS results with the pristine NbSe2, in which there is a VHS at -0.2 eV with respect to the Fermi level, particularly due to saddle points in the electronic band structure near the M point in the Brillouin zoneCalandra et al. (2009). The intercalated transition metal d-orbitals bands are narrower than the Nb bands and closer to the Fermi level. Once the Nb and Ni bands hybridize, the Nb bands are pushed away from the Fermi level. The intercalation of the Ni atoms shifts the VHS to -0.4 eV from the Fermi level, as we can see in both Figs. 13(a) and 13(b). As a result, this shift of the VHS reduces the TDOS at the Fermi level and can suppress the electronic instabilities, such as the CDW and superconductivityYan et al. (2019). It was shown that there is a strong interplay between a VHS and a CDW state in the isostructural and isoelectronic 2H-TaSe2Luckin et al. (2024). Therefore, we can expect that Ni intercalation would affect the CDW. Indeed, after intercalation, the change of the Fermi surface helps stabilize the lattice. On the other hand, Ni electrons are involved in the Nb-Se covalent bonds within each monolayer, which weakens the Nb-Nb interactions and prevents the clustering effect of Nb atoms responsible for the CDW distortion. Thus, because of these two factors, the CDW phase at low temperature is expected to be suppressed in the intercalated NbSe2.Yin et al. (2023).
Another property of this family is the large orbital magnetic moment. At U=2 eV for the FM and AFM2 phase, the spin moment is around 0.9 . In these phases, the orbital magnetic moment is parallel to the spin-moment and it is 0.09 and 0.08 for the FM and AFM2, respectively. Therefore, the orbital magnetic moment is of the order of 10% with respect to the spin-moment.
III.4 Non-collinear phases due to magnetic exchanges
@@
The magnetic coupling is antiferromagnetic within the plane; therefore, our system is an effective 2D triangular lattice with antiferromagnetic first-neighbor interaction for low U and ferromagnetic first-neighbor interaction for high U. The collinear antiferromagnetic phase investigated in the previous section is the stripe phase. From the literature, we know that the ground state of a 2D triangular system is the stripe when the second neighbor magnetic exchange is antiferromagnetic above a threshold which is around =0.14; while the ground state is non-collinear 120∘ phase when the second neighbor interaction is ferromagnetic or weakly antiferromagneticHu et al. (2015). Between these two phases, the theory predicts the quantum spin liquid phase for insulating systems. Since the system is metallic, it cannot host the quantum spin liquid phase, but the competition between the stripe and the non-collinear phases holds. We constructed a supercell to simulate the non-collinear 120∘ phase, which is reported in Fig. 14. We performed the structural relaxation on the non-collinear phase in the presence of SOC, and we obtained that the non-collinear phase is 15 meV per formula unit higher in energy than the stripe. Therefore, we can assume that the value of is antiferromagnetic and it is the key parameter to stabilize the stripe AFM2 phase. Another factor that assisted the magnetic coupling in stabilizing the stripe phase is the large magnetocrystalline anisotropy. Indeed, the stripe phase can have spins along the easy axis (z-axis), while the non-collinear 120∘ phase has in-plane spins.
III.5 Weak ferromagnetism in the altermagnetic phase of T0.25MX2
We study the weak ferromagnetism in the altermagnetic phases of TMX2 and T0.25MX2 (T = 3d, M = 4d/5d, X= S, Se, Te). What is relevant for the weak ferromagnetism calculations is only the crystal and magnetic symmetries. Since the symmetries are the same for all the T0.25MX2 systems, their properties such as weak ferromagnetism would be the same. The only quantity that is material dependent is the magnitude of the net magnetic moment; thus, this calculation depends on the symmetries of the magnetic space group and the main results are valid for all intercalated TMDs with this stoichiometry. The weak ferromagnetism is a characteristic of systems with zero magnetization which break time-reversal symmetry, where the spin-orbit coupling generates an antisymmetric exchange interaction, leading to spin canting and a net magnetization. The simplest mechanism that produces spin canting is the staggered Dzyaloshinskii–MoriyaAutieri et al. (2025), but higher-order antisymmetric magnetic exchanges were also reported depending on the symmetries of the systemRoig et al. (2025).
Defining the x-axis as the direction between the magnetic atoms and the y-axis as the orthogonal direction in the ab-plane, we consider the Néel vector N orientations along the x-, y-, and the z-axis. The altermagnetic phase contains two magnetic atoms that we define as Ni1 and Ni2, which have magnetic moments M1 and M2, respectively. The Néel vector here is defined as the difference between the magnetizations of the two magnetic sublattices N=M1 - M2 of an altermagnet. In the case of TMX2, when the Néel vector is along the x-axis or z-axis, we have no weak ferromagnetism. When the Néel vector is along the y-axis, we have weak ferromagnetism along the z-axis. This is the same case as the NiAs structure compound, such as MnTe and CrSb.
In the case of T0.25MX2 (T = 3d, M = 4d/5d, X=S, Se, Te), although the crystal space group remains the same, the presence of more atoms reduces the symmetries of the magnetic space group. While the Te atoms in MnTe occupy the Wyckoff position 2c, in Ni0.25NbSe2, there are none in this position. With more atoms as Nb(1), Nb(2), Se(1) and Se(2), T0.25MX2 is expected to have lower symmetry. Therefore, we expect more cases with weak ferromagnetism. We are not aware of any prior studies addressing weak ferromagnetism in intercalated compounds with this space group. In this case, we always have weak ferromagnetism for the Néel vector along z and y, while we have weak ferrimagnetism for the Néel vector along z. Therefore, this is different from the NiAs structure, and the anomalous Hall effect (AHE) is always allowed in this class of materials.
The results of the spin cantings due to the antisymmetric exchanges in altermagnets are reported in Fig. 15. We report the spin-density, but from the properties of the spin-density, we can infer the symmetry property of the spin-components. When the Néel vector is along the x-axis, the relativistic spin-density of Sx for Ni1 and Ni2 are equal and opposite; therefore, we have S=-S as reported in Fig. 15(a). The other components are around 3 orders of magnitude lower; we have S=S and S=-S as shown in Fig. 15(b) and (c), respectively. Therefore, we have weak ferromagnetism along the y-axis and a small rotation of the Néel vector along the z-axis. When the Néel vector is along the y-axis, the relativistic spin density of Sy for Ni1 and Ni2 are equal and opposite, so we have S=-S as demonstrated in Fig. 15(e). The other two spin components Sx and Sz are 4 and 3 orders of magnitude smaller than Sy respectively, where we have S=-S and S=S as shown in Fig. 15(d) and (f). We can deduce that we have weak ferromagnetism along the z-axis and a small rotation of the Néel vector along the y-axis. Finally, placing the Néel vector along the z-axis, the relativistic spin density of Sz for Ni1 and Ni2 are equal and opposite as can be seen in Fig. 15(i). Similarly, the other two components are three orders of magnitude smaller than Sz. This case is, however, different than the other two cases, where the other two components are neither equal nor opposite. In this case, we have weak ferrimagnetism present along the x-axis and the z-axis.
The summary of the weak ferromagnetism and the weak ferrimagnetism is reported in Fig. 16. In all cases, the main components are equal and opposite; therefore, the weak ferromagnetism in this material class is observed only orthogonal to the Néel vector and the spins are three orders of magnitude smaller than the main components. Regarding the AHE (which has the same symmetries as the weak ferromagnetism), we can observe from Fig. 16(a,b) that 0 for Nx, while 0 for Ny. In the last case with Nz, we have 0 and 0 as shown in Fig. 16(c,d). In the previous subsection, the noncollinearity originated from the frustration of the symmetric magnetic exchanges, while the phases studied in this subsection are non-collinear due to the antisymmetric exchange interactions.
While in the literature orbital ferromagnetism in altermagnets was highlightedYe et al. (2025); Jo et al. (2025), this class of altermagnets shows orbital antiferromagnetism. Tuning the Néel vector, we obtain three inequivalent cases with large orbital moments in directions orthogonal to the spins when the Néel vector is in the ab-plane and parallel to the spins when the Néel vector is along the z-axis. When the Néel vector is along the x-axis, we have weak ferromagnetism along the y-axis with the spin magnetic moment of the order of 0.001 ; however, the orbital moment is 0.07 along the y-axis. For the Néel vector along the y-axis, we have weak ferromagnetism along the z-axis; however, the orbital moments are along the x-axis with the size of 0.07 . Finally, when the Néel vector is along the z-axis, the orbital magnetic moment is 0.1 parallel to the spin moment. As found in the literature for other altermagnets, the orbital magnetic moment is 1-2 orders of magnitude larger than the weak ferromagnetismYe et al. (2025); Jo et al. (2025). Also in this case, the orbital moment is around 10% of the total spin moment.
@@ IV Conclusions
We performed first-principle calculations for the Ni-intercalated NbSe2 focusing on the stoichiometry Ni0.25NbSe2. We observe a strong stability of the intercalated Ni atoms to be in the middle between two Nb atoms with different z-coordinate but the same x- and y-coordinates. Our results indicate that the magnetic properties of this system are highly tunable and capable of hosting a rich variety of magnetic phases.
@@
Ni0.25NbSe2 is at the border between a ferromagnetic state and a stripe-like antiferromagnetic phase with Kramers’ degeneracy (referred to as the AFM2 phase). This configuration is characterized by strong out-of-plane ferromagnetic coupling and relatively weak in-plane magnetic interactions. Among the magnetic structures considered, the AFM2 phase emerges as the ground state in the regime of low Coulomb repulsion, consistent with experimental conditions, where screening effects and a deviation from half-filling reduces the effective interaction strength. The system exhibits an easy axis along and a strong in-plane second-nearest-neighbor antiferromagnetic interaction. We calculate the electric properties of NixNbSe2 (x=0.25, ,and 1) and found that the Ni connectivity strongly influences the ground state. When the Ni atoms belonging to two different layers are one on top of each other, the magnetic coupling between the layers is ferromagnetic. When the Ni atoms belonging to two different layers are shifted, the magnetic coupling between the layers is antiferromagnetic and the ground state can be altermagnetic. Ni intercalation significantly alters the electronic structure near the Fermi level and makes the electronic properties more three-dimensional with respect to the pristine NbSe2. It shifts the VHS away from the Fermi energy, reducing the TDOS and suppressing electronic instabilities. Nevertheless, the system remains metallic, with a substantial DOS at the Fermi level that provides effective screening of Coulomb interactions. Notable changes are observed in the Fermi surface compared to pristine NbSe2, with multiple sheets appearing in both the FM and AFM2 phases. In the AFM2 and FM phases, there is a switch in the type of charge carriers at the point where an electron pocket appears at dominated by 3z2-r2 orbital character, with the primary atomic weight localized on the Nb atoms.
Finally, we report the study of the weak ferromagnetism for Ni0.25NbSe2. This can serve as benchmark for the family members of T0.25MX2 (T = 3d, M = 4d/5d, X= S, Se, Te). We find weak ferromagnetism for an in-plane Néel vector, while weak ferrimagnetism appears for the out-of-plane Néel vector. This differs from the NiAs structure, which has the same space group but not the same magnetic space group. We conclude that the family of intercalated T0.25MX2 hosts AHE for every orientation of the Néel vector. Additionally, we find a large orbital magnetic moment with antiferromagnetic order, which differs from previously studied altermagnets where orbital ferromagnetism was found.
We consider what may occur as we move to more complex cases. In the case of disordered Ni distribution or deviation of the Ni content from 1/3 and 1/4, the fact that Ni intercalated NbS2 has an antiferromagnetic coupling on the triangular lattice in the ab-plane and out-of-plane interactions likely drives experimental samples into a spin-glass regime or magnetically disordered system. Given the geometrical frustration and proximity to both ferromagnetic and antiferromagnetic regimes, a metamagnetic transition under an applied magnetic field is also plausible. Assuming a proportionality between the Ni concentration and Néel temperature, the critical temperature is expected to be below 84 K, which is the Néel temperature of Ni1/3NbS2An et al. (2023). We look forward to this magnetic system’s experimental realization, which offers a promising platform for exploring tunable magnetism and correlated electronic behavior.